OpenStudy (anonymous):

Simplify if possible. If an expression cannot be simplified, write "not possible" as the answer. 3√2-2√3

OpenStudy (ghazi):

\[3=\sqrt{3}\sqrt{3}\] and \[2= \sqrt{2}\sqrt{2}\] can you simplify it now?

OpenStudy (anonymous):

No, I'm still really confused.

OpenStudy (ghazi):

\[\sqrt{3}\sqrt{3}\sqrt{2}-\sqrt{2}\sqrt{2}\sqrt{3}= \]

OpenStudy (ghazi):

now?

OpenStudy (anonymous):

Oh god.

OpenStudy (anonymous):

I have no idea what is going on in this problem.

OpenStudy (ghazi):

\[\sqrt{3}*\sqrt{2}(\sqrt{3}-\sqrt{2})\]

OpenStudy (anonymous):

do I multiply 3 with the square root of 2?

OpenStudy (ghazi):

yes

OpenStudy (ghazi):

\[\sqrt{6}(\sqrt{3}-\sqrt{2})\]

OpenStudy (ghazi):

is it clear now?

OpenStudy (anonymous):

No, I'm still really lost.

OpenStudy (ghazi):

i have just broken 3 =\[\sqrt{3}*\sqrt{3}\] similarly 2 and then took common from them..check out the above comments

OpenStudy (anonymous):

I just need to mainly know the order of how to solve the problem. I don't want it rearranged or reworded, I just wanna know if it should look something like this... 3(√2)-2(√3)

OpenStudy (anonymous):

√2= 1.41, correct?

OpenStudy (ghazi):

yes that is what i have done....i have simplified it further....but it is fine too...and yes that's correct

OpenStudy (anonymous):

ok, so then 3 multiplied by the √2 would be like... 4.24?

OpenStudy (ghazi):

yes..

OpenStudy (anonymous):

so would I take (4.24)-(2(√3))?

OpenStudy (anonymous):

4.24-3.46?

OpenStudy (ghazi):

you can also take \[2\sqrt{3}=2*1.732=3.46\]

OpenStudy (ghazi):

yes,now subtract and find the exact value

OpenStudy (anonymous):

so the answer would be .78?

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

Changing irrationals into decimals is generally frowned upon by math instructors. What you started with is about as simple as it gets. You can change it a little, but not make it "simpler". I think your answer was "not possible".

OpenStudy (anonymous):

well it is possible, it's just frowned upon, yes?

OpenStudy (anonymous):

Actually, it is not possible to represent an irrational number as a decimal accurately. You must round.

OpenStudy (anonymous):

to the nearest 10th, or what?

OpenStudy (anonymous):

Rounding makes it wrong.

OpenStudy (anonymous):

I don't wanna put down not possible though and then they get upset that I didn't try and simplify it.

OpenStudy (anonymous):

is it breaking any mathematical rules or laws to turn these into decimals?

OpenStudy (anonymous):

Ghazi had great ideas for manipulating it... but I think that is your answer.

OpenStudy (ghazi):

well it's better to leave it under square roots....but it's fine if you have to reduce it further

OpenStudy (anonymous):

because they ask me for the square root of 3, and 5, and 6 multiple times in future questions.

OpenStudy (anonymous):

Do they give you advice on how to round for those... like "to the nearest..."

OpenStudy (anonymous):

No, they don't tell me anything, which is why I'm a little flustered about it.

OpenStudy (anonymous):

Good point.

OpenStudy (anonymous):

they do have an example... √0.09= 0.3

OpenStudy (anonymous):

A math teacher usually wants exact... meaning with the radicals. An engineer or scientist would round, so they can use it in real life.

OpenStudy (anonymous):

Your example is exact.

OpenStudy (anonymous):

so should I round to the nearest tenth... or?

OpenStudy (anonymous):

.8?

OpenStudy (anonymous):

I would mark it wrong. If I were your instructor, I would be looking for "not possible".

OpenStudy (anonymous):

ok, I guess that's what I'll put